# Complete transparency with three active-passive-coupled optical resonators

### v1

@@ Appendix A Appendix
The system dynamics can be as well understood when there is an obvious gain saturation effect. For any possible gain saturation intensity $I_{S}$, the gain rate $g_{2}$ after saturation (even in the processes of dynamically unstable regimes) will become a constant, so that the system can be nonetheless described by a group of linear differential equations for its coupled intracavity modes. The difference is that the saturated gain rate, $g_{2,s}=\kappa_{2,0}/\left(1+\frac{<\hat{a}^{\dagger}_{2}\hat{a}_{2}>_{s}}{I_{S}}\right)$ ($\hat{a}^{\dagger}_{2}\hat{a}_{2}>_{s}$ is the stabilized photon number in the active resonator), will be lowered from the original $\kappa_{2,0}$, and the complete transparency point will be moved to another coupling value $J_{2}=\sqrt{(g_{2,s}-\gamma_{2})\kappa_{3}}$ as long as the saturated effective gain rate, $g_{2,s}-\gamma_{2}$, is still positive. In the dynamically stable regimes determined by the original gain rate $\kappa_{2,0}$ (the effective gain rate $\kappa_{2}$ in Fig. S-2 is replaced by another constant $\kappa_{2,0}-\gamma_{2}$), the intracavity field intensity or the photon number $\hat{a}^{\dagger}_{2}\hat{a}_{2}>$ in the active cavity will cause a rather limited shift of the gain rate $g_{2}$ from the original $\kappa_{2,0}$, so that the complete transparency can be achieved by a relatively fine tuning of the coupling $J_{2}$ to a smaller value; see an example in Appendix A. It is why the complete transparency can be always reached at a coupling $J_{2}$ in such dynamically stable regimes, given any feasible gain saturation intensity $I_{S}$. One feature is that tuning the coupling $J_{2}$ around the point of complete transparency will not lead to a drastic change of the photon number $\hat{a}^{\dagger}_{2}\hat{a}_{2}>$ in the active resonator. Especially, at the exact point of complete transparency, there is a counter-intuitive photon number $\hat{a}^{\dagger}_{2}\hat{a}_{2}>_{s}=2\kappa_{ex}(\varepsilon_{p}/J_{1})^{2}$ [from Eq. (S-12)] that is independent of the coupling $J_{2}$ and the gain rate, and it provides a mechanism to reduce the gain saturation effect by an enhanced coupling $J_{1}$ [exhibited by the results in Fig. S-1(c)]. On the other hand, the optical gain will drop tremendously due to a quick growth of the cavity photon number $\hat{a}^{\dagger}_{2}\hat{a}_{2}>$ in a dynamically unstable regime, and the stabilized gain rate $g_{2,s}$ will be far away from the original $\kappa_{2,0}$, making it impossible to reach the complete transparency point by tuning the coupling $J_{2}$.
@@
In summary, the phenomenon of complete transparency through the concerned structure of three active-passive-coupled optical resonators should be observed in the dynamically stable regimes determined by its built-in gain rate $\kappa_{2,0}$, which can become smaller with a gain saturation. The system’s dynamical stability can be guaranteed if it exists with this original gain rate $\kappa_{2,0}$, because the effective gain will be only lowered after the saturation. With a negligible gain saturation effect, i.e., the field intensity in the active resonator is not high enough to modify the gain rate obviously, the complete transparency point will be located by adjusting the distance between cavity $2$ and $3$ so that the corresponding coupling goes across the horizontal line $J_{2}=\sqrt{\kappa_{2}\kappa_{3}}$ in Fig. S-2(a), and the operation can be performed with any nonzero coupling $J_{1}$ between cavity $1$ and cavity $2$. Even if there exists an obvious gain saturation that shifts the stabilized optical gain rate, the point of complete transparency can still be located by a further tuning of the coupling $J_{2}$ to a bit smaller value. Moreover, the gain saturation effect can be well suppressed by enhancing the coupling strength $J_{1}$.

### v2

@@ Appendix A Appendix
The system dynamics can be as well understood when there is an obvious gain saturation effect. For any possible gain saturation intensity $I_{S}$, the gain rate $g_{2}$ after saturation (even in the processes of dynamically unstable regimes) will become a lowered constant, so that the system can be nonetheless described by a group of linear differential equations for its coupled intracavity modes. The saturated gain rate, $g_{2,s}=\kappa_{2,0}/\left(1+\frac{<\hat{a}^{\dagger}_{2}\hat{a}_{2}>_{s}}{I_{S}}\right)$ ($\hat{a}^{\dagger}_{2}\hat{a}_{2}>_{s}$ is the stabilized photon number in the active resonator), will be lowered from the original $\kappa_{2,0}$, and the complete transparency point will be moved to another coupling value $J_{2}=\sqrt{(g_{2,s}-\gamma_{2})\kappa_{3}}$, as long as the saturated net gain rate, $g_{2,s}-\gamma_{2}$, is still positive. In the dynamically stable regimes determined by the original gain rate $\kappa_{2,0}$, i.e., the net gain rate $\kappa_{2}$ in Fig. S-2 is replaced by another constant $\kappa_{2,0}-\gamma_{2}$, the intracavity field intensity or the photon number $\hat{a}^{\dagger}_{2}\hat{a}_{2}>$ in the active cavity will cause a rather limited shift of the gain rate $g_{2}$ from the original $\kappa_{2,0}$, so that the complete transparency can be achieved by a relatively fine tuning of the coupling $J_{2}$ to a smaller value; see an example in Sec. I. It is why the complete transparency can be always reached at a coupling $J_{2}$ in such dynamically stable regimes, given any feasible gain saturation intensity $I_{S}$. An interesting feature is that tuning the coupling $J_{2}$ around the point of complete transparency will not lead to a drastic change of the photon number $\hat{a}^{\dagger}_{2}\hat{a}_{2}>$ in the active resonator. Especially, at the exact point of complete transparency, there is a counter-intuitive photon number $\hat{a}^{\dagger}_{2}\hat{a}_{2}>_{s}=2\kappa_{ex}(\varepsilon_{p}/J_{1})^{2}$ [from Eq. (S-12)] that is independent of the coupling $J_{2}$ and the net gain rate, and it provides a mechanism to reduce the gain saturation effect by an enhanced coupling $J_{1}$ [exhibited by the results in Fig. S-1(c)]. On the other hand, the optical gain will drop tremendously due to a quick growth of the cavity photon number $\hat{a}^{\dagger}_{2}\hat{a}_{2}>$ in the dynamically unstable regime, and the stabilized gain rate $g_{2,s}$ will be thus far away from the original $\kappa_{2,0}$, making it impossible to reach the complete transparency point by tuning the coupling $J_{2}$.
@@
In summary, the phenomenon of complete transparency through the concerned structure of three active-passive-coupled optical resonators should be observed in the dynamically stable regimes determined by its built-in gain rate $\kappa_{2,0}$, which can become smaller with a gain saturation. The system’s dynamical stability can be guaranteed if it exists with this original gain rate $\kappa_{2,0}$, because the net gain rate will be only lowered after its saturation so that the system becomes more stable. With a negligible gain saturation effect, i.e., the field intensity in the active resonator is not high enough to modify the gain rate obviously, the complete transparency point will be located by adjusting the distance between cavity $2$ and $3$ so that the corresponding coupling goes across the horizontal line $J_{2}=\sqrt{\kappa_{2}\kappa_{3}}$ in Fig. S-2(a), and the operation can be performed with any nonzero coupling $J_{1}$ between cavity $1$ and cavity $2$. Even if there exists an obvious gain saturation that shifts the stabilized optical gain rate, the point of complete transparency can still be located by a further tuning of the coupling $J_{2}$ to a bit smaller value. Moreover, the gain saturation effect can be well suppressed by enhancing the coupling strength $J_{1}$.